Circular No. 8487

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

V2361 CYGNI = NOVA CYGNI 2005 N. N. Samus, Institute of Astronomy, Moscow, inform us that the

designation V2361 Cyg has been given to this nova (cf. IAUC 8483, 8484).

S. Ciroi, F. Di Mille, and P. Rafanelli, Department of Astronomy, University of Padova, report that a spectrum (range 446–680 nm; typical resolution 1.5 nm) of this nova, taken on Feb. 12.195 UT at the 1.2-m telescope of the Asiago Astrophysical Observatory (+ Boller & Chivens spectrograph) is dominated by the Balmer emission lines Hα and Hβ with at least three absorption systems on their blue wings, characterized by velocities of −3200, −6500, and −7500 km/s. Emission lines of Fe II multiplets 37, 38, 42, 49, 57, and 74; of O I 9 and 10; and of N II 3, 36, 63, and 66 are present in the spectrum.

CCD V magnitudes forwarded by E. O. Waagen, AAVSO: Feb. 12.4637 UT, 10.13 (S. Dvorak, Clermont, FL); 13.7750, 10.28 (R. Miles, Stourton Caundle, Dorset, England); 14.7985, 10.40 (Miles). Visual magnitude estimates: Feb. 14.183, 10.2 (R. J. Bouma, Groningen, The Netherlands); 15.138, 10.3 (K. Hornoch, Lelekovice, Czech Republic).

SUPERNOVA 2005ai IN NGC 2314 S. Taubenberger, Max-Planck-Institut für Astrophysik, Garching; F.

Patat, European Southern Observatory; and S. Benetti, Istituto Nazionale di Astrofisica, Padua, on behalf of the ERTN (cf. IAUC 7987), report that a fully reduced spectrum (range 360–880 nm; resolution 1.3 nm) of SN 2005ai (cf. IAUC 8486), obtained on Feb. 14.83 UT by M. Alises at the Calar Alto Observatory 2.2-m telescope (+ CAFOS), shows it to be a type-Ia supernova. The spectrum is very similar to that of SN 1994D (Patat et al. 1996, MNRAS 278, 111) at ∼ 1 month past maximum light.

COMET C/1997 E2 (SOHO)

An additional Kreutz sungrazing comet (cf. IAUC 8483):

Comet 1997 UT α_{2000} δ_{2000} Inst. F MPEC
C/1997 E2 Mar. 5.281 23^{h}22.^{m}2 − 6^{o}57′ C3 XL 2005-C38

SUPERNOVA 2005af IN NGC 4945 Visual magnitude estimates: Feb. 14.826 UT, 13.3 (A. Pearce, Ned-

lands, W. Australia); 15.135, 12.6 (W. Souza, São Paulo, Brazil).

2005 February 15  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green